OpenStudy (anonymous):

In triangle ABC, we are given AB + BC = 34 and BC + AC = 34. Classify triangle ABC Scalene Isosceles Equilateral Right .

OpenStudy (anonymous):

Isosceles

OpenStudy (anonymous):

AB + BC = 34 = BC + AC AB + BC = BC + AC AB + BC - BC = BC - BC + AC AB = AC Therefore 2 sides are equal, however it is impossible to tell whether BC is equal to AB or AC. e.g., BC could equal 5 while AC and AB both equal 29, or BC could equal to 17 while AC and AB are both 17 as well. So it must be Isosceles

OpenStudy (anonymous):

isosceles...

OpenStudy (anonymous):

thank you, y'all are amazing (: